The perimeter and the area of a rectangle are 48 cm and 140 cm square. Find the length and the width of the rectangle
Answer by mananth(16946)

perimeter = 48
let length = l
width =w
2(l+w)=48
l+w=24................(1)
area = l*w
l*w=140
l=140/w
substitute l in (1)
140/w + w = 24
multiply equation by w
140+w^2=24w
w^2-24w+140=0
w^2-14w-10w+140=0
w(w-14)-10(w-14)=0
(w-14)(w-10)=0
w=10 OR 14
dimensions are 14cm and 10cm